Suppose the average grade in this class is 85 with standard deviation 2.50.

a) If each grade is increased by 5 points, what are the new mean and standard deviation?

b) If each grade is increased by 5%, what are the new mean and standard deviation?
Answer #1

a)Mean will increase by 5 and standard deviation will remain same , or mean = 90, standard deviation = 2.5

b) Mean will increase by 5% and standard deviaion will also increase by 5% .

Mean = 85*1.05=89.25

Standard deviation = 2.5*1.05 = 2.625